If your are travelling at 3.5km/h, 30.0 degrees north of east what is you northerly component of your velocity?


1
Expert's answer
2021-03-21T11:39:38-0400
"v_y=v_N=v_0sin\theta=3.5\ \dfrac{km}{h}\cdot sin30^{\circ}=1.75\ \dfrac{km}{h}."